Write a polynomial equation with the roots
a) 2,5,1
Can you please help me out? Thanks so much in advance:)
Can you also please show the steps it would really help me understand:)

Answer by jim_thompson5910(35256)

x = 2, x = 5 or x = 1

x - 2 = 0, x - 5 = 0 or x - 1 = 0

(x - 2)(x - 5)(x - 1) = 0

(x - 2)(x^2 - 6x + 5) = 0

x(x^2 - 6x + 5)-2(x^2 - 6x + 5) = 0

x^3 - 6x^2 + 5x-2x^2 + 12x - 10 = 0

x^3 - 8x^2 + 17x - 10 = 0

Final Answer: y = x^3 - 8x^2 + 17x - 10